OpenStudy (anonymous):

Graph the solution to the following inequality on the number line. (x+2)(x-7)>0

hero (hero):

(x + 2)(x - 7) >0 x^2 - 5x - 14 > 0 Graph x^2 - 5x - 14

hero (hero):

Then observe the areas above the x axis where the graph is true.

OpenStudy (anonymous):

Okay, how did you get rid of the parenthesis?

hero (hero):

I multiplied it out. You know how to multiply polynomials, right?

hero (hero):

(x + 2)(x - 7) = x(x - 7) + 2(x - 7) = x^2 - 7x + 2x - 14 = x^2 - 5x - 14

OpenStudy (anonymous):

Ohh, right right.

hero (hero):

(x + 2)(x - 7) = x(x - 7) + 2(x - 7) = x^2 - 7x + 2x - 14 = x^2 - 5x - 14

hero (hero):

Do you know what the solution is now?

OpenStudy (anonymous):

I'd forgotten about that. So how do I get x^2 - 5x - 14 onto a number line? Just starting school again and am still getting into the swing of things.

OpenStudy (anonymous):

It makes sense for a graph, but number line?

hero (hero):

I will try to explain it to you.

hero (hero):

Do you at least remember how to find the zeroes of the quadratic expression?

OpenStudy (anonymous):

Yes

hero (hero):

What are the zeroes?

hero (hero):

Okay. I guess I'll just explain

OpenStudy (anonymous):

To get the zeros wouldn't we have to take it back to binomials?

hero (hero):

The zeroes of the graph are x = -2 , x = 7 Because x^2 -5x - 14 factors to (x + 2)(x - 7) and using the zero product property x + 2 = 0 x = -2 x - 7 = 0 x = 7

hero (hero):

We graph that on the number line. 

OpenStudy (anonymous):

Got it! I guess I was just overthinking the problem

hero (hero):

We're not done yet...

hero (hero):

Trust me..there's more

hero (hero):

Allow me to finish completely

OpenStudy (anonymous):

Okay

hero (hero):

Next we find the vertex of the parabola since a quadratic expression is graphed as a parabola... We use the vertex of a parabola formula to find the vertex: \[x = -\frac{b}{2a}\] a = 1 b = -5 \[x = \frac{5}{2} = 2 \frac{1}{2}\]

hero (hero):

then we have to insert that into x to find y: We need to graph the quadratic expression: After inputting x and solving for y we get: y = (5/2)^2 -5(5/2) - 14 = -41/4 or -10 1/4

hero (hero):